OpenStudy (anonymous):

Solve for y m=1/2(x+y)

OpenStudy (anonymous):

\[solve for y : m=\frac{ 1 }{ 2 }(x+y)\]

OpenStudy (anonymous):

multiply both sides by 2 get \[2m=x+y\] subract \(x\) get \[2m-x=y\]